OpenStudy (anonymous):

simplify square root of 80 - the square root of 45 + the square root of 500

OpenStudy (anonymous):

\[\sqrt{80}-\sqrt{45}+\sqrt{500}=\sqrt{16*5}-\sqrt{9*5}+\sqrt{100*5}\] \[4\sqrt{5}-3\sqrt{5}+10\sqrt{5}=11\sqrt{5}\]